OpenStudy (anonymous):

how do i state the domain for these functions? 1. f(x) = 3x / x^2-5 2. h(x)= x+2 / √x^2-7

OpenStudy (anonymous):

\[\frac{ 3x }{ x^2-5 }\]?

OpenStudy (anonymous):

\[h(x)=\frac{ x+2 }{ \sqrt{x^2-7} }\]?

OpenStudy (zzr0ck3r):

@wagnermar use more parenthesis please, as you would in a calculator.

OpenStudy (anonymous):

@Johnbc yes thats correct. @zzr0ck3r sorry

OpenStudy (goformit100):

Yes it's correct

OpenStudy (anonymous):

?

OpenStudy (anonymous):

what is correct?

OpenStudy (goformit100):

The posted replies are correct.

OpenStudy (anonymous):

yes, it's correct.

OpenStudy (anonymous):

do you know how to solve these problems? i need help

OpenStudy (goformit100):

Sorry ma'm. I am really sorry. Please excuse me.

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

the domains of the functions are the values of the independant variable for which the functions are defined. can you tell me when the first function is not defined.

OpenStudy (anonymous):

x^2-5 ?

OpenStudy (anonymous):

Here's a hint; an expression is undefined when you divide by zero.

OpenStudy (anonymous):

x^2-5=0?

OpenStudy (anonymous):

Exactly. Now you need to solve for x. This tells us what x-value the function is not defined for. Because when x is this soon-to-be-found value, you are dividing by zero, which is undefined.

OpenStudy (anonymous):

x^2=5

OpenStudy (anonymous):

Yupp.... keep on going. You haven't solved for x yet. You are done when you have\[x=some-expression\]

OpenStudy (anonymous):

x=√5 and x=-√5 ?

OpenStudy (anonymous):

Correct. Therefore, the first function is defined whenever x is not \[\pm \sqrt{5}\]That means the domain would be all real numbers except plus or minus square root of 5.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

The other function is not defined when the denominator is zero, but it also has another limitation. Do you know what that is?

OpenStudy (anonymous):

??? ≠ ??? i dont know?

OpenStudy (anonymous):

What is the square root of a negative number?

OpenStudy (anonymous):

Trick question. The square root of a negative number is undefined.

OpenStudy (anonymous):

a square root of a negative number isnt possible right?

OpenStudy (anonymous):

Correct. Therefore, the function is only defined when x squared minus seven is greater than zero.

OpenStudy (anonymous):

ok